A dB meter registers 123 dB when placed 3.0 m in front of a loudspeaker. Assuming uniform spherical spreading of the sound, what is the power output of the speaker? IN W
Answer #1

Sound level (in dB) = 10 log 10(I/I 0 ) where I is intensity,

I 0 = 10 -12 W/ m 2

So , according to problem, 10 log (I/ I 0 )= 123

=> log (I/I 0 ) = 12.3

=> I/ I 0 = 10 12.3 => I = 10-12 * 10 12.3 = 10 0.3 W/m 2

Power= intensity * area, area= 4* pi* r 2 , where r is radius

So, power= 10 0.3 *4*pi*3*3 = 225.66 W